OpenStudy (anonymous):

Is the line through points P(–8, –10) and Q(–5, –12) perpendicular to the line through points R(9, –6) and S(17, –5)? Explain.

hero (hero):

no, because when the lines intersect they should intersect at 90 degree angles. They do not.

OpenStudy (anonymous):

\[\tan \theta=\frac{m_1-m_2}{1+m_1m_2}\] \[m=\frac{y_2-y_1}{x_2-x_1}\] \[m_1=\frac{-5-(-6)}{17-9}=1/8\] \[m_2=\frac{-12-(-10)}{-5-(-8)}=-2/3\] \[\tan \theta=\frac{-2/3-1/8}{1+(-2/3)*1/8)}=19/22\] \[\theta=\tan^{-1} (19/22)\] \[\theta=40.8\]

hero (hero):

talk about taking it further than necessary

OpenStudy (anonymous):